# Global Continuation of Stable Periodic Orbits in Systems of Competing Predators

### v1

@@ 2 Proof of analytic families of periodic orbits
$x_{j}(t):=\frac{1}{\kappa}\frac{m_{j}}{\gamma y_{j}}X_{j}\left(\frac{1}{\gamma}\tau t\right),\qquad s(t):=\frac{1}{\kappa}S\left(\frac{1}{\gamma}\tau t\right),$ (5)
$\alpha_{j}:=\frac{1}{\kappa}a_{j},\qquad\lambda_{j}:=\frac{1}{\kappa}\left(\frac{a_{j}d_{j}}{m_{j}-d_{j}}\right),\qquad\delta_{j}:=\frac{m_{j}-d_{j}}{\gamma},$
@@
$\eta\in[-1,1]\quad\mapsto\quad\kappa(\eta):=\frac{2\kappa_{1}\kappa_{2}}{\kappa_{1}+\kappa_{2}+(\kappa_{1}-\kappa_{2})\eta}$ (7)
@@
In the next sections, we set up an auxiliary system for which the family of periodic orbits is locally isolated, so that no transcritical bifurcations occur at the boundary planes $Q_{1}$ and $Q_{2}$. Then, we formulate a zero-finding problem whose zeros are in one-to-one correspondence with families of periodic solutions to the rescaled system (6). Next, we derive a set of sufficient conditions for a Newton-type fixed-point operator to be a contraction near a high-order numerical approximation of the family.
2.1 Desingularization
@@
To isolate the branch, we use a blow-up (as in “zoom-in”) method. This strategy is inspired by [11, 16, 50]. We consider $u=(u_{1},u_{2},u_{3})$ with
@@
The new variables $\zeta_{1}$ and $\zeta_{2}$ correspond to amplitude-like parameters. As we impose $u_{j}(0,\eta)=1$ for $j=1,2$, then $\zeta_{j}(\eta)=x_{j}(0,\eta)$. This leads to the following auxiliary system for the continuation parameter $\eta\in[-1,1]$:
@@
Given $\zeta_{1}$ and $\zeta_{2}$, we have that $u=(u_{1},u_{2},u_{3})$ is a periodic solution to the auxiliary system (10) if and only if $(\zeta_{1}u_{1},\zeta_{2}u_{2},u_{3})$ is a periodic solution to the rescaled system (6). Provided that $\zeta_{1},\zeta_{2}\neq 0$, then the periodic orbit is positive. Importantly, periodic orbits are locally isolated whenever $\zeta_{j}=0$. For completeness, we state this in the following lemma. However, we stress that verifying its hypotheses is unnecessary for our proof; this step merely serves to justify that (10) provides a suitable system for our contraction argument (detailed in Section 2.3) to work.
Lemma 2.1.
Let $j,j^{\prime}\in\{1,2\}$ with $j^{\prime}\neq j$. Suppose that there exists a boundary limit cycle of the rescaled system (6) in $Q_{j^{\prime}}$ for some $\eta\in[-1,1]$. If $\zeta_{j}=0$, then the auxiliary system (10) has an isolated solution $u=(u_{1},u_{2},u_{3})$ with $(u_{j^{\prime}},u_{3})$ the periodic solution to
$\begin{cases}\displaystyle\partial_{t}u_{j^{\prime}}=\tau\delta_{j^{\prime}}\left(\frac{u_{3}-\lambda_{j^{\prime}}(\eta)}{u_{3}+\alpha_{j^{\prime}}(\eta)}\right)u_{j^{\prime}},\\ \displaystyle\partial_{t}u_{3}=\tau\left(1-u_{3}-\zeta_{j^{\prime}}\frac{u_{j^{\prime}}}{u_{3}+\alpha_{j^{\prime}}(\eta)}\right)u_{3}.\end{cases}$ (11)
Proof.
Note that the equations (11) satisfied by $(u_{j^{\prime}},u_{3})$ are those of (10) when $\zeta_{j}=0$, and for $u_{j}$ we have the linear differential equation
$\partial_{t}u_{j}=\tau\delta_{j}\left(\frac{u_{3}-\lambda_{j}}{u_{3}+\alpha_{j}}\right)u_{j}.$ (12)
Now, since $Q_{j^{\prime}}$ admits a boundary limit cycle, it must be that
$\int_{0}^{2\pi}\frac{u_{3}(t,\eta)-\lambda_{j}(\eta)}{u_{3}(t,\eta)+\alpha_{j}(\eta)}\,\mathrm{d}t=0.$
Thus, the scalar linear equation (12) admits a family of periodic solutions, parameterized by its amplitude, whose phase is determined by $u_{3}$. Hence, the linear scaling $u_{j}(0,\eta)=1$ characterizes a locally unique periodic orbit. ∎
Furthermore, we must fix the phase of the periodic solutions to remove their time-translation invariance. To do so, we follow [15] and impose the phase condition
@@
[…]
2.2 Zero-finding problem
We now formulate a system of functional equations, whose zero corresponds to a family of periodic solutions to the auxiliary system (10). We denote the $n$-th Chebyshev coefficient of a function $\psi\in C^{\infty}([-1,1],\mathbb{C})$, namely
$\psi_{n}=\frac{1}{2\pi}\int_{-1}^{1}\frac{\psi(\eta)\mathcal{T}_{n}(\eta)}{\sqrt{1-\eta^{2}}}\,\mathrm{d}\eta,\qquad n\in\mathbb{N}_{0}:=\mathbb{N}\cup\{0\}.$ (14)
Note that with this convention
$\psi(\eta)=\sum_{n\in\mathbb{Z}}\psi_{|n|}\mathcal{T}_{|n|}(\eta)=\psi_{0}+2\sum_{n\geq 1}\psi_{n}\mathcal{T}_{n}(\eta).$
Here $\mathcal{T}_{n}:[-1,1]\mapsto[-1,1]$ represents the $n$-th Chebyshev polynomial of the first kind, given by the recurrence relation
$\mathcal{T}_{0}(\eta)=1,\qquad\mathcal{T}_{1}(\eta)=\eta,\qquad\mathcal{T}_{n}=2\eta\mathcal{T}_{n-1}(\eta)-\mathcal{T}_{n-2}(\eta),\qquad n\geq 2.$ (15)
Importantly, these polynomials satisfy the identity
$\mathcal{T}_{n}(\cos(\theta))=\cos(n\theta),$ (16)
meaning that an infinite series of Chebyshev polynomials amounts to a cosine series. This feature highlights the significance of employing them as a basis with respect to $\eta$, since the convergence property of analytic Fourier series holds true for functions defined on the entire range of parameter values $[-1,1]$. This fact contrasts with Taylor expansions, where despite the analyticity of a function, the presence of poles often necessitates partitioning its domain into several subintervals.
For $\nu\geq 1$, we define
$\mathcal{P}_{\nu}:=\left\{\psi\in C^{\infty}([-1,1],\mathbb{C})\,:\,\|\psi\|_{\mathcal{P}_{\nu}}<\infty\right\},$ (17)
where, given $\psi\in C^{\infty}([-1,1],\mathbb{C})$,
$\|\psi\|_{\mathcal{P}_{\nu}}:=\sum_{n\in\mathbb{Z}}|\psi_{|n|}|\nu^{|n|}=\frac{1}{2\pi}\sum_{n\in\mathbb{Z}}\left|\int_{-1}^{1}\frac{\psi(\eta)\mathcal{T}_{|n|}(\eta)}{\sqrt{1-\eta^{2}}}\,\mathrm{d}\eta\right|\nu^{|n|}.$ (18)
The choice of $\nu\geq 1$ is related to the regularity of the function $\psi$. For $\nu=1$, $\mathcal{P}_{\nu}$ amounts to the Wiener algebra. For $\nu>1$, $\psi$ is analytic inside the Bernstein ellipse $\left\{z\in\mathbb{C}\,:\,|z|\leq\frac{1}{2}|\nu e^{i\theta}+\nu^{-1}e^{-i\theta}|,\,\theta\in(-\pi,\pi]\right\}$ and term-by-term differentiation is well defined up to any order, which will play an important role to verify that our family of periodic orbits crosses each of the boundary planes $Q_{1}$ and $Q_{2}$ exactly once; see Section 2.4.
Lemma 2.2.
$\mathcal{P}_{\nu}$ is a unital Banach algebra with respect to the multiplication of functions, specifically
$\|\phi\psi\|_{\mathcal{P}_{\nu}}\leq\|\phi\|_{\mathcal{P}_{\nu}}\|\psi\|_{\mathcal{P}_{\nu}},\qquad\text{for all }\phi,\psi\in\mathcal{P}_{\nu},$ (19)
where
Proof.
We have
$\|\phi\psi\|_{\mathcal{P}_{\nu}}$ $=\sum_{n\in\mathbb{Z}}\Big{|}\sum_{n^{\prime}\in\mathbb{Z}}\phi_{|n-n^{\prime}|}\psi_{|n^{\prime}|}\Big{|}\nu^{|n|}$
$\leq\sum_{n\in\mathbb{Z}}\sum_{n^{\prime}\in\mathbb{Z}}|\phi_{|n-n^{\prime}|}||\psi_{|n^{\prime}|}|\nu^{|n^{\prime}|}\nu^{|n|-|n^{\prime}|}$
$\leq\left(\sum_{n\in\mathbb{Z}}|\phi_{|n|}|\nu^{|n|}\right)\left(\sum_{n^{\prime}\in\mathbb{Z}}|\psi_{|n^{\prime}|}|\nu^{|n^{\prime}|}\right)$
$=\|\phi\|_{\mathcal{P}_{\nu}}\|\psi\|_{\mathcal{P}_{\nu}}.$
@@ 2.2 Zero-finding problem
$\phi_{n,k}=\frac{1}{(2\pi)^{2}}\int_{0}^{2\pi}\int_{-1}^{1}\frac{\phi(t,\eta)e^{ikt}\mathcal{T}_{n}(\eta)}{\sqrt{1-\eta^{2}}}\,\mathrm{d}\eta\,\mathrm{d}t,\qquad n\in\mathbb{N}_{0},\quad k\in\mathbb{Z}.$ (22)
For $\nu\geq 1$, we define
$\mathcal{W}_{\nu}:=\left\{\phi\in C^{\infty}(\mathbb{R}/2\pi\mathbb{Z}\times[-1,1],\mathbb{C})\,:\,\|\phi\|_{\nu}<\infty\right\}$ (23)
where, given $\phi\in C^{\infty}(\mathbb{R}/2\pi\mathbb{Z}\times[-1,1],\mathbb{R})$,
$\|\phi\|_{\nu}:=\sum_{n,k\in\mathbb{Z}}|\phi_{|n|,k}|\nu^{|n|}=\frac{1}{(2\pi)^{2}}\sum_{n,k\in\mathbb{Z}}\left|\int_{0}^{2\pi}\int_{-1}^{1}\frac{\phi(t,\eta)e^{ikt}\mathcal{T}_{|n|}(\eta)}{\sqrt{1-\eta^{2}}}\,\mathrm{d}\eta\,\mathrm{d}t\right|\nu^{|n|}.$ (24)
Lemma 2.3.
$\mathcal{W}_{\nu}$ is a unital Banach algebra with respect to the multiplication of functions, specifically
@@
Proof.
The proof is similar to the one of Lemma 2.2. ∎
Our objective is to show that a fixed-point operator $T$ (yet to be constructed) is a contraction around a finite Fourier–Chebyshev series that approximates the family of periodic orbits. This approximation lives in a finite-dimensional subspace of the Banach space
$\mathcal{X}_{\nu}:=\mathcal{P}_{\nu}\times\mathcal{P}_{\nu}\times\mathcal{P}_{\nu}\times\mathcal{U}_{\nu},$ (27)
endowed with the norm
$\|\chi\|_{\mathcal{X}_{\nu}}:=\|\tau\|_{\mathcal{P}_{\nu}}+\|\zeta_{1}\|_{\mathcal{P}_{\nu}}+\|\zeta_{2}\|_{\mathcal{P}_{\nu}}+\|u\|_{\mathcal{U}_{\nu}},\qquad\text{for all }\chi=(\tau,\zeta_{1},\zeta_{2},u)\in\mathcal{X}_{\nu},$ (28)
where
$\mathcal{U}_{\nu}:=\left\{u=(u_{1},u_{2},u_{3})\in\mathcal{W}_{\nu}^{3}\,:\,\|u\|_{\mathcal{U}_{\nu}}:=\sum_{j=1}^{3}\|u_{j}\|_{\nu}<\infty\right\}.$ (29)
To obtain the finite-dimensional subspace of $\mathcal{X}_{\nu}$, we introduce the truncation operators $\Pi_{K},\Pi_{N,K}:C^{\infty}(\mathbb{R}/2\pi\mathbb{Z}\times[-1,1],\mathbb{C})\to C^{\infty}(\mathbb{R}/2\pi\mathbb{Z}\times[-1,1],\mathbb{C})$ by
$(\Pi_{K}u)_{k}:=\begin{cases}u_{k},&|k|\leq K,\\ 0,&|k|>K,\end{cases},\qquad(\Pi_{N,K}u)_{k}:=\begin{cases}(\hat{\Pi}_{N}u_{k})_{n},&|k|\leq K,\\ 0,&|k|>K,\end{cases}$ (30)
@@
$(\hat{\Pi}_{N}u)_{n}:=\begin{cases}u_{n},&n\leq N,\\ 0,&n>N.\end{cases}$ (31)
@@
Fix $N,K\in\mathbb{N}$ and $\partial_{t}\Gamma\in\Pi_{N,K}\mathcal{U}_{\nu}$. A family of periodic orbits to the auxiliary system (10), parameterized by $\eta\in[-1,1]$, is a zero of the nonlinear unbounded operator $F:\mathcal{D}(F)\subset\mathcal{X}_{\nu}\to\mathcal{X}_{\nu}$ defined, for all $\chi=(\tau,\zeta_{1},\zeta_{2},u)\in\mathcal{X}_{\nu}$, $u=(u_{1},u_{2},u_{3})$, by
$F(\chi):=\begin{pmatrix}\rho(u)\\ \partial_{t}u-\tau f(\zeta_{1},\zeta_{2},u)\end{pmatrix},$ (34)
where
$\rho(u)$ $:=\begin{pmatrix}\displaystyle\int_{0}^{2\pi}\langle u(t,\,\cdot),\partial_{t}\Gamma(t,\,\cdot)\rangle\,\mathrm{d}t\\ [\Pi_{N,K}u_{1}](0,\,\cdot)-1\\ [\Pi_{N,K}u_{2}](0,\,\cdot)-1\end{pmatrix},$ (35)
$f(\zeta_{1},\zeta_{2},u)$ $:=\begin{pmatrix}\displaystyle\delta_{1}\left(\frac{u_{3}-\lambda_{1}}{u_{3}+\alpha_{1}}\right)u_{1}\\ \displaystyle\delta_{2}\left(\frac{u_{3}-\lambda_{2}}{u_{3}+\alpha_{2}}\right)u_{2}\\ \displaystyle\left(1-u_{3}-\zeta_{1}\frac{u_{1}}{u_{3}+\alpha_{1}}-\zeta_{2}\frac{u_{2}}{u_{3}+\alpha_{2}}\right)u_{3}\end{pmatrix}.$ (36)
We emphasize again that $\alpha_{j}=\alpha_{j}(\eta),\lambda_{j}=\lambda_{j}(\eta)$ are parameterized by $\eta$ as described in (8). The domain of $F$ is given by
2.3 Verifying the contraction
We begin by finding an approximation of the branch of periodic orbits in $\Pi_{N,K}\mathcal{X}_{\nu}$, denoted by $\bar{\chi}=(\bar{\tau},\bar{\zeta}_{1},\bar{\zeta}_{2},\bar{u})$. For $\ell=0,\dots,N$, we consider the Chebyshev node
$\eta_{\ell}:=\cos\left(\frac{\ell}{N}\pi\right).$ (38)
Denoting $F|_{\eta_{\ell}}$ the mapping $F$ given by (34) evaluated at $\eta=\eta_{\ell}$, we solve numerically $N+1$ problems of finite dimension by using Newton’s method on $\Pi_{0,K}\circ F|_{\eta_{\ell}}\circ\Pi_{0,K}$, and we denote by $\bar{\chi}_{\ell}\in\Pi_{0,K}\mathcal{X}_{\nu}$ the approximate zero, that is, $\Pi_{0,K}F|_{\eta_{\ell}}(\bar{\chi}_{\ell})\approx 0$. Then, we retrieve numerically a Chebyshev polynomial associated to each component of $\bar{\chi}_{\ell}$ via the inverse discrete Fourier transform, thereby yielding $\bar{\chi}\in\Pi_{N,K}\mathcal{X}_{\nu}$. Incidentally, it is practical to opt for a suitable number of Chebyshev nodes to take the full advantage of the fast Fourier transform algorithm.
Given two Banach spaces $\mathfrak{X},\mathfrak{Y}$, we denote by $\mathscr{B}(\mathfrak{X},\mathfrak{Y})$ the set of bounded linear operators from $\mathfrak{X}$ to $\mathfrak{Y}$. To prove the existence of a zero of $F$ defined in (34), we consider the fixed-point operator
$T(\chi):=\chi-AF(\chi),$ (39)
where $A\in\mathscr{B}(\mathcal{X}_{\nu},\mathcal{X}_{\nu})$ is given by
$A:=A_{\textnormal{finite}}\Pi_{K}+A_{\textnormal{tail}}\Pi_{>K},$ (40)
with $A_{\textnormal{finite}}\in\mathscr{B}(\Pi_{K}\mathcal{X}_{\nu},\Pi_{K}\mathcal{X}_{\nu})$ an approximation of the inverse $(\Pi_{N,K}DF(\bar{\chi})\Pi_{N,K})^{-1}$ and $A_{\textnormal{tail}}:\Pi_{>K}\mathcal{X}_{\nu}\to\Pi_{>K}\mathcal{X}_{\nu}$ is defined, for all $\chi=(\tau,\zeta_{1},\zeta_{2},u)\in\mathcal{X}_{\nu}$, $u=(u_{1},u_{2},u_{3})$, by

### v2

@@ 2 Proof of analytic families of periodic orbits
$x_{j}(t)\coloneqq\frac{1}{\kappa}\frac{m_{j}}{\gamma y_{j}}X_{j}\left(\frac{1}{\gamma}\tau t\right),\qquad s(t)\coloneqq\frac{1}{\kappa}S\left(\frac{1}{\gamma}\tau t\right),$ (16)
$\alpha_{j}\coloneqq\frac{1}{\kappa}a_{j},\qquad\lambda_{j}\coloneqq\frac{1}{\kappa}\left(\frac{a_{j}d_{j}}{m_{j}-d_{j}}\right),\qquad\delta_{j}\coloneqq\frac{m_{j}-d_{j}}{\gamma},$
@@
$\eta\in[-1,1]\quad\mapsto\quad\kappa(\eta)\coloneqq\frac{2\kappa_{1}\kappa_{2}}{\kappa_{1}+\kappa_{2}+(\kappa_{1}-\kappa_{2})\eta}$ (18)
@@
The roadmap for proving the existence of the family is the following:
  • Formulation: we set up a zero-finding problem whose zeros are in one-to-one correspondence with families of periodic solutions to the rescaled system (17).
  • Analysis: we derive a set of sufficient conditions for a Newton type fixed-point operator to be a contraction near a high-order numerical approximation of the family.
  • Computations: we perform the desired computations to verify the contraction conditions; see the code available in [32]. The code incorporates some relevant checks discussed in Section 2.3 to ensure that the proven family is real-valued and indeed crosses the boundary planes $Q_{1}$ and $Q_{2}$. It also contains the proof of the stability, which is based on a contraction argument as well, though for a different set of equations to solve.
2.1 Zero-finding problem
@@
Therefore, this falls in the scope of Section 1.2, and we consider $u=(u_{1},u_{2},u_{3})$ with
@@
This leads to the following auxiliary system for the continuation parameter $\eta\in[-1,1]$:
@@
Then, a family of periodic orbits to the auxiliary system (21), parameterized by $\eta\in[-1,1]$, is a zero of the map $F$ defined (at this stage only formally), for $\chi=(\tau,\zeta_{1},\zeta_{2},u)$, by
$F(\chi)\coloneqq\begin{pmatrix}\rho(u)\\ \partial_{t}u-\tau f(\zeta_{1},\zeta_{2},u)\end{pmatrix}$, (22)
where, $u=(u_{1},u_{2},u_{3})$,
$\rho(u)$ $\coloneqq\begin{pmatrix}\displaystyle\int_{0}^{2\pi}\langle u(t,\,\cdot),\partial_{t}\Gamma(t,\,\cdot)\rangle\,\mathrm{d}t\\ \ell(u_{1})-1\\ \ell(u_{2})-1\end{pmatrix}$, (23)
$f(\zeta_{1},\zeta_{2},u)$ $\coloneqq\begin{pmatrix}\displaystyle\delta_{1}\left(\frac{u_{3}-\lambda_{1}}{u_{3}+\alpha_{1}}\right)u_{1}\\ \displaystyle\delta_{2}\left(\frac{u_{3}-\lambda_{2}}{u_{3}+\alpha_{2}}\right)u_{2}\\ \displaystyle\left(1-u_{3}-\zeta_{1}\frac{u_{1}}{u_{3}+\alpha_{1}}-\zeta_{2}\frac{u_{2}}{u_{3}+\alpha_{2}}\right)u_{3}\end{pmatrix}$, (24)
for some map $\ell$ mimicking the normalization constraints $u_{1}(0,\eta)=u_{2}(0,\eta)=0$, and given explicitly at the end of this section. We emphasize again that $\alpha_{j}=\alpha_{j}(\eta),\lambda_{j}=\lambda_{j}(\eta)$ are parameterized by $\eta$ as described in (19). The constraint
@@
where $\langle\cdot,\cdot\rangle$ denotes the standard inner product on $\mathbb{C}^{3}$, fixes the phase of the periodic solutions to remove their time-translation invariance. Here, $\Gamma$ denotes some periodic function near the one we seek; its time derivative $\partial_{t}\Gamma$ may simply be estimated numerically and described by a finite number of Fourier–Chebyshev coefficients as specified at the end of this section.
We now detail the functional analytic framework used to define $F$. The approach is spectral. We denote the $n$-th Chebyshev coefficient of a function $\psi\in C^{\infty}([-1,1],\mathbb{C})$, namely
$\psi_{n}=\frac{1}{2\pi}\int_{-1}^{1}\frac{\psi(\eta)T_{n}(\eta)}{\sqrt{1-\eta^{2}}}\,\mathrm{d}\eta,\qquad n\in\mathbb{N}_{0}\coloneqq\mathbb{N}\cup\{0\}.$ (26)
Here $T_{n}:[-1,1]\mapsto[-1,1]$ represents the $n$-th Chebyshev polynomial of the first kind, satisfying the identity
$T_{n}(\cos(\theta))=\cos(n\theta).$ (27)
@@
$\phi_{n,k}=\frac{1}{4\pi^{2}}\int_{0}^{2\pi}\int_{-1}^{1}\frac{\phi(t,\eta)e^{ikt}T_{n}(\eta)}{\sqrt{1-\eta^{2}}}\,\mathrm{d}\eta\,\mathrm{d}t,\qquad n\in\mathbb{N}_{0},\quad k\in\mathbb{Z}.$ (29)
For $\nu\geq 1$, we introduce
$\mathcal{P}_{\nu}$ $\coloneqq\left\{\psi\in C^{\infty}([-1,1],\mathbb{C})\,:\,\|\psi\|_{\mathcal{P}_{\nu}}\coloneqq\sum_{n\in\mathbb{Z}}|\psi_{|n|}|\nu^{|n|}<\infty\right\},$ (30)
$\mathcal{W}_{\nu}$ $\coloneqq\left\{\phi\in C^{\infty}(\mathbb{R}/2\pi\mathbb{Z}\times[-1,1],\mathbb{C})\,:\,\|\phi\|_{\nu}\coloneqq\sum_{k\in\mathbb{Z}}\|\phi_{k}\|_{\mathcal{P}_{\nu}}=\sum_{n,k\in\mathbb{Z}}|\phi_{|n|,k}|\nu^{|n|}<\infty\right\}.$ (31)
The choice of $\nu\geq 1$ is related to the regularity of the function $\psi$; see [62] for an extensive discussion. For our purposes, we note that, whenever $\nu>1$, term-by-term differentiation of the series is well defined up to any order, which will play an important role to verify that our family of periodic orbits crosses each of the boundary planes $Q_{1}$ and $Q_{2}$ exactly once; see Section 2.3. Unlike Taylor series, even non-analytic functions may be represented by Chebyshev series, and in the analytic case, we always have convergence of the series in a Bernstein ellipse.
Furthermore, $\mathcal{P}_{\nu}$ forms a Banach algebra with respect to the multiplication of functions, specifically
[…]
where, for all $\phi,\psi\in\mathcal{P}_{\nu}$, we have
Similarly, $\mathcal{W}_{\nu}$ forms a Banach algebra
@@
where, for all $\phi,\psi\in\mathcal{W}_{\nu}$, we have
Consider the Banach space
$\mathcal{X}_{\nu}\coloneqq\mathcal{P}_{\nu}\times\mathcal{P}_{\nu}\times\mathcal{P}_{\nu}\times\mathcal{U}_{\nu},$ (36)
endowed with the norm
$\|\chi\|_{\mathcal{X}_{\nu}}\coloneqq\|\tau\|_{\mathcal{P}_{\nu}}+\|\zeta_{1}\|_{\mathcal{P}_{\nu}}+\|\zeta_{2}\|_{\mathcal{P}_{\nu}}+\|u\|_{\mathcal{U}_{\nu}},\qquad\text{for all }\chi=(\tau,\zeta_{1},\zeta_{2},u)\in\mathcal{X}_{\nu},$ (37)
@@
$\mathcal{U}_{\nu}\coloneqq\left\{u=(u_{1},u_{2},u_{3})\in\mathcal{W}_{\nu}^{3}\,:\,\|u\|_{\mathcal{U}_{\nu}}\coloneqq\sum_{j=1}^{3}\|u_{j}\|_{\nu}<\infty\right\}.$ (38)
Then, we have $F:\mathcal{D}(F)\subset\mathcal{X}_{\nu}\to\mathcal{X}_{\nu}$ with
Our strategy is to show that a fixed-point operator (yet to be constructed) is a contraction around a finite Fourier–Chebyshev series that approximates the family of periodic orbits. This approximation lives in a finite-dimensional subspace of $\mathcal{X}_{\nu}$. To obtain this space, we introduce the truncation operators $\Pi_{K},\Pi_{N,K}:C^{\infty}(\mathbb{R}/2\pi\mathbb{Z}\times[-1,1],\mathbb{C})\to C^{\infty}(\mathbb{R}/2\pi\mathbb{Z}\times[-1,1],\mathbb{C})$ by
$(\Pi_{K}u)_{k}\coloneqq\begin{cases}u_{k},&|k|\leq K,\\ 0,&|k|>K,\end{cases},\qquad(\Pi_{N,K}u)_{k}\coloneqq\begin{cases}(\hat{\Pi}_{N}u_{k})_{n},&|k|\leq K,\\ 0,&|k|>K,\end{cases}$ (40)
@@
$(\hat{\Pi}_{N}u)_{n}\coloneqq\begin{cases}u_{n},&n\leq N,\\ 0,&n>N.\end{cases}$ (41)
@@
Fix $N,K\in\mathbb{N}$. We consider
$\partial_{t}\Gamma\in\Pi_{N,K}\mathcal{U}_{\nu},\qquad\ell(\psi)=[\Pi_{N,K}\psi](0,\,\cdot),\quad\text{for all }\psi\in\mathcal{W}_{\nu}.$ (44)
With this choice of $\partial_{t}\Gamma$ and $\ell$, the map $\rho$ given in(23) only depends on a finite number of Fourier–Chebyshev coefficients.
2.2 Newton–Kantorovich type theorem
The next theorem provides sufficient conditions for $T$ to be a contraction which, by construction of $A$, is enough to obtain a zero of $F$. The proof is standard and can be found, for instance, in [13].
Theorem 2.1.
Let $R>0$, $\bar{\chi}\in\Pi_{N,K}\mathcal{X}_{\nu}$, and $A:\mathcal{X}_{\nu}\to\mathcal{X}_{\nu}$ be an injective linear map. Suppose there exist positive constants $Y,Z_{1},Z_{2}=Z_{2}(R)$ satisfying
$\|AF(\bar{\chi})\|_{\mathcal{X}_{\nu}}$ $\leq Y,$ (45a)
$\|ADF(\bar{\chi})-I\|_{\mathscr{B}(\mathcal{X}_{\nu},\mathcal{X}_{\nu})}$ $\leq Z_{1},$ (45b)
$\sup_{\chi\in B_{R}(\bar{\chi})}\|AD^{2}F(\chi)\|_{\mathscr{B}(\mathcal{X}_{\nu},\mathscr{B}(\mathcal{X}_{\nu},\mathcal{X}_{\nu}))}$ $\leq Z_{2},$ (45c)
where $B_{R}(\bar{\chi})$ is the closed ball in $\mathcal{X}_{\nu}$ centered at $\bar{\chi}$ with radius $R$, and $\mathscr{B}(\mathcal{X},\mathcal{Y})$ denotes the set of bounded linear operators from $\mathcal{X}$ to $\mathcal{Y}$.
If
$2YZ_{2}$ $\leq(1-Z_{1})^{2},$ (46a)
$Z_{1}$ $<1,$ (46b)
then, for any $r>0$ such that
$\frac{1-Z_{1}-\sqrt{(1-Z_{1})^{2}-2YZ_{2}}}{Z_{2}}\leq r<\min\left(\frac{1-Z_{1}}{Z_{2}},R\right),$ (47)
there exists a unique zero $\tilde{\chi}\in B_{r}(\bar{\chi})$ of $F$.
Before verifying the contraction criteria given in the above theorem, we must first have an approximation of the branch of periodic orbits in $\Pi_{N,K}\mathcal{X}_{\nu}$, denoted by $\bar{\chi}=(\bar{\tau},\bar{\zeta}_{1},\bar{\zeta}_{2},\bar{u})$. This can be achieved by interpolating each Fourier mode of the approximate zeros $\bar{\chi}_{0},\dots,\bar{\chi}_{N}$ of $F|_{\eta_{\ell}}$, where $F|_{\eta_{\ell}}$ is the mapping defined in (22) and evaluated at $\eta=\eta_{\ell}$. A practical way is to consider a finite truncation of $F|_{\eta_{\ell}}$ and use Newton’s method to produce an accurate $\bar{\chi}_{\ell}$ at each of the $N+1$ Chebyshev nodes
$\eta_{\ell}\coloneqq-\cos\left(\frac{\ell\pi}{N}\right),\qquad\ell=0,\dots,N.$ (48)
Then, the corresponding Fourier coefficients can be recovered by applying the inverse discrete Fourier transform (recall identity (27)). In practice, this step can be carried out efficiently by means of fast Fourier transform algorithms.
Moreover, we need to construct the operator $A\in\mathscr{B}(\mathcal{X}_{\nu},\mathcal{X}_{\nu})$. We set
$A\coloneqq A_{\textnormal{finite}}\Pi_{K}+A_{\textnormal{tail}}\Pi_{>K},$ (49)
with $A_{\textnormal{tail}}:\Pi_{>K}\mathcal{X}_{\nu}\to\Pi_{>K}\mathcal{X}_{\nu}$ defined, for all $\chi=(\tau,\zeta_{1},\zeta_{2},u)\in\mathcal{X}_{\nu}$, $u=(u_{1},u_{2},u_{3})$, by